Solve the system of equations using the Addition/Elimination Method.

x – 3y = 3
0.5x + 2y = –0.25

Found 3 solutions by Tobiasz, Alan3354, richwmiller:
Answer by Tobiasz(54)

-0.5(x – 3y = 3 )
0.5x + 2y = –0.25
-0.5x - 1.5y = 1.5 Should be -0.5x + 1.5y = 1.5
0.5 + 2y = -0.25
0.5y = 1.25 <---- divide both by 0.5
y+=+2.5 <--- plug this in any equation and you get x
x - 2.5 = 3
x+=+0.5
(0.5,2.5)
----------
Check:
x – 3y = 3
0.5 - 3*2.5 =? 3
-7 =? 3 No.
--------------
0.5x + 2y = –0.25
0.5*0.5 + 2*2.5 =? -0.25
.25 + 5 =? -0.25 No
---------------------
So it's wrong.


Answer by Alan3354(69443)

x – 3y = 3
0.5x + 2y = –0.25 --> -x - 4y = 0.5
------------------
-x -4y = 0.5
x – 3y = 3
---------------- Add
-7y = 3.5
y = -0.5
----------
x = 1.5

Answer by richwmiller(17219)

x=1.5, y~-0.5
x – 3y = 3
0.5x + 2y = –0.25
multiply second by 2
x-3y=3
x+4y=-.5
subtract second from first
-7y=3.5
y=-.5
x-3*-.5=3
x+1.5=3
x=1.5